A man of mass 10kg is standing on a plank so that he is 20m from a point 'A'(point 'A' is not on plank).He walks 8m on the plank towards 'A' and stops.The mass of the plank is 40kg and frictipn between plank and ground is negligible.How far is the man from point 'A' now? 

Grade:11

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Sindhuji P,

Ans:- Let the distance travelled by the man on the plank is = x mt, mass of the man= m kg and the plank=M kg

then the plank will move horizontally in the opposite direction a distance y.

So, the net displacement of the man is (x-y)

hence the eq is,   m(x-y)=My

y=mx/(m+M)

given that x=8 m

hence y=1.6 mt. So the net displacement of the man is=(8-1.6)= 6.4 mt

Hence the present distance from point A is =(20-6.4)=13.6 mt (ans)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Sindhuji P !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur